OpenStudy (anonymous):

Find the unit vector that has the same direction as the vector v. v = 3i + j

OpenStudy (anonymous):

the unit vector with the same direction as v can be found it v divided by its magnitude

OpenStudy (anonymous):

So it should be v/abs(v) (3i + j)*(1/abs(v)) (3i + j)/(3^2+1^2)^(1/2) 3i/(10)^(1/2) + j/(10)^(1/2) or (3i + j)/(10)^(1/2)

OpenStudy (anonymous):

A. u =3/sq rt of 10 i + 1/sq rt of 10 j B. u = 3/sq rt of 11 i + 1/sq rt of 11 j C. u = 3sq rt of 10 i + sq rt of 10 j D. u = 10/sq rt of 3 i + sq rt of 10 j

OpenStudy (anonymous):

\[unit V= \frac{ 3i+j }{ \sqrt {3^2+1^2} }=\frac{ (3i+j) }{ \sqrt{10} }\]

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

;) keep rocking